For what values of k does 3x^2 - kx + 2 = 0 have equal roots
Can you please help me out thanks so much in advance:)

Answer by stanbon(75887)

For what values of k does 3x^2 - kx + 2 = 0 have equal roots
----
A quadratic has equal roots when the discriminant = 0
----
b^2 - 4ac = 0
k^2 - 4*3*2 = 0
k^2 = 24
k = 2sqrt(6) or k = -2sqrt(6)
----------------
Cheers,
Stan H.